### v1

# Uniqueness of sign-changing solutions to Trudinger’s equation

Abstract.
We establish uniqueness for sign-changing solutions to Trudinger’s parabolic equation with time dependent $C^{2}$ Dirichlet boundary data.
This work was supported by the Research Council of Finland (project 360185), and the Finnish Ministry of Education and Culture’s Pilot for Doctoral Programmes (Pilot project Mathematics of Sensing, Imaging and Modelling). We thank Verena Bögelein for providing a reference.
1. Introduction
We consider the doubly non-linear evolutionary equation of Trudinger [Tru68], which reads
$\frac{\partial}{\partial t}\left(\left|u\right|^{p-2}u\right)\,=\,\operatorname{div}(\left|\nabla u\right|^{p-2}\nabla u),\quad 1<p<\infty,$
in the bounded domain $\Omega_{T}=\Omega\times(0,T),$ where $\Omega\subset{\mathbb{R}}^{n}$ is a Lipschitz domain. In [Tru68] N. Trudinger observed that, due to the homogeneous structure, no ’intrinsic scaling’ is needed in Harnack’s inequality. This was later proved by J. Kinnunen and T. Kuusi [KK07] as well as by U. Gianazza and V. Vespri [GV06]. In particular, when it comes to sign-changing solutions, little is known about Trudinger’s equation in contrast to the situation for the evolutionary $p$-Laplace equation
$\frac{\partial u}{\partial t}\,=\,\operatorname{div}(|\nabla u|^{p-2}\nabla u).$
Especially, under natural assumptions the uniqueness of sign-changing solutions is a challenging open problem for Trudinger’s equation. For solutions $u$ with $\inf\{u\}>0$ this is settled, see [LL22] for references.
We shall focus on the uniqueness of sign-changing weak solutions belonging to the Sobolev space $L^{p}(0,T;W^{1,p}(\Omega))$. Our main result, Theorem 4.1, establishes uniqueness for sign-changing solutions with $C^{2}$ boundary data. For solutions with time-independent Dirichlet boundary values on the lateral boundary $\partial\Omega\times[0,T]$ F. Otto proved uniqueness in [Ott96]: for example weak solutions in $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ are unique. In fact, these solutions turn out to have a time derivative $\partial_{t}(|u|^{p-2}u)$. Less regular solutions in $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ cannot exist! We extend Otto’s result to time dependent boundary data, yet of class $C^{2}$. The idea of our proof is, first following Otto’s approach in [Ott96], to use Kružkov’s [Kru70] doubling of time variables to establish the comparison principle, but now extended to time dependent boundary values. The argument requires some space between the boundary values. Thus an extra device is needed: we finish the proof by considering the auxiliary solutions
$\begin{cases}u_{+\varepsilon}\quad\text{with boundary values}\quad\psi+{\varepsilon}\\ u_{-\varepsilon}\quad\text{with boundary values}\quad\psi-{\varepsilon}.\end{cases}$
An arbitrary weak solution $u$ with boundary values $\psi$ will be squeezed
$u_{-\varepsilon}\,\leq\,u\,\leq\,u_{+\varepsilon}.$
We can conclude the proof by using the $C^{2}$ assumption as in [LPS25a] to prove that both auxiliary solutions converge to the same solution as $\varepsilon\to 0$.
It has not escaped our attention that the $C^{2}$ assumption can be somewhat relaxed. Attempts to prove Theorem 3.1 for $\varepsilon=0$ have hitherto not succeeded.
2. Preliminaries
Throughout the paper, $\Omega\subseteq{\mathbb{R}}^{n}$ is a bounded Lipschitz domain, and we consider the fixed parameters $1<p<\infty$ and $0<T<\infty$. We use the standard notation $\Omega_{T}:=\Omega\times(0,T)$,
$\partial_{p}\Omega_{T}:=(\overline{\Omega}\times\{0\})\cup(\partial\Omega\times[0,T]),$
and denote the parabolic Sobolev space by $L^{p}(0,T;W^{1,p}(\Omega))$. Recall that this function space consists of measurable functions $u:\Omega_{T}\to{\mathbb{R}}$ such that the function $x\mapsto u(x,t)$ belongs to $W^{1,p}(\Omega)$ for almost every $t\in(0,T)$, and $u$ has the finite norm
$\biggl(\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\left|u\right|^{p}+\left|\nabla u\right|^{p}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\biggr)^{\frac{1}{p}}<\infty.$
Here $\nabla u$ denotes the spatial gradient. The function space $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ is defined analogously. We use the function $b$ defined by
$b(z):=\left|z\right|^{p-2}z,\qquad b(0)=0$
for real values. We write $z^{+}=\max\{z,0\}$.
Definition 2.1.
Let $\psi\in C(\overline{\Omega_{T}})\cap L^{p}(0,T;W^{1,p}(\Omega))$, and
$u\in L^{p}(0,T,W^{1,p}(\Omega))$. We say that $u$ is a weak solution with the (Dirichlet) boundary values $\psi$ if the following conditions hold.
  1. (1)
    $u-\psi\in L^{p}(0,T;W_{0}^{1,p}(\Omega))$.
  2. (2)
    For every $\gamma\in C_{0}^{\infty}(\Omega_{T})$ we have
[…]
  3. (3)
    $u$ satisfies the initial value condition
    (2.2) $\lim_{t\to 0+}\int\limits_{\Omega}\left|b(u(x,t))-b(\psi(x,t))\right|\hskip 0.50003pt\textup{d}x\,=\,0.$
Given $\psi$ as above, a weak solution $u$ always exists with the given initial condition, see [AL83] and also[MN23]. By [BDL21] weak solutions are locally Hölder continuous. By Theorem 1.2 and Theorem 1.3 in [BDL21], the boundary values are continuously attained in $\partial\Omega\times(0,T)$ and $\Omega\times\{0\}$.
3. Comparison principle
In this section a suitable comparison principle for the uniqueness proof is provided. Our method is strongly influenced by Otto’s work [Ott96], but includes some necessary modifications to handle time-dependent boundary values. The method originates from the work of S. Kružkov [Kru70]. See [Ott96, Section 4] for further discussion.
It will be important not to assume any existence of time derivatives. Indeed, under such an extra assumption the result is well-known. For the convenience of the reader, we present a standard proof following the notation in [LPS25a]: Consider two solutions $u_{1},u_{2}$ that have time derivatives $\partial_{t}(|u_{j}|^{p-2}u_{j})$ at least in $L^{1}(\Omega_{T})$ and the same boundary values. Let $H_{\delta}(s)$ denote the approximation
$H_{\delta}(s)\,=\,\begin{cases}0,\,\,\,s\leq 0\\ \dfrac{s}{\delta},\,\,\,0<s<\delta\\ 1,\,\,\,s\geq\delta.\end{cases}$
of the Heaviside function. We use the test function $\zeta=H_{\delta}\bigl(u_{2}-u_{1}\bigr)$ in the equation
$\int\limits_{t_{1}}^{t_{2}}\!\!\int\limits_{\Omega}\Bigl(\zeta\,\partial_{t}(|u_{j}|^{p-2}u_{j})\,+\,\big\langle|\nabla u_{j}|^{p-2}\nabla u_{j},\nabla\zeta\big\rangle\Bigr)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\,=\,0,\quad j=1,2.$
Subtracting the equations, we get
$\int\limits_{t_{1}}^{t_{2}}\!\!\int\limits_{\Omega}\partial_{t}\left(|u_{2}(x,t)|^{p-2}u_{2}(x,t)-|u_{1}(x,t)|^{p-2}u_{1}(x,t)\right)H_{\delta}\bigl(u_{2}-u_{1}\bigr)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t$
$=-\frac{1}{\delta}\int\!\!\int\!\big\langle|\nabla u_{2}|^{p-2}\nabla u_{2}-|\nabla u_{1}|^{p-2}\nabla u_{1},\nabla u_{2}-\nabla u_{1}\big\rangle\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\,\,\leq\,\,0,$
by the vector inequality $\big\langle|b|^{p-2}b-|a|^{p-2}a,b-a\big\rangle\,\geq\,0$. Letting $\delta\to 0$ we obtain
$\int\limits_{t_{1}}^{t_{2}}\!\!\int\limits_{\Omega}\partial_{t}\left(|u_{2}(x,t)|^{p-2}u_{2}(x,t)-|u_{1}(x,t)|^{p-2}u_{1}(x,t)\right)^{+}\!\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\,\,\leq\,\,0.$
This implies
$\int\limits_{\Omega}\left(|u_{2}(x,t)|^{p-2}u_{2}(x,t)-|u_{1}(x,t)|^{p-2}u_{1}(x,t)\right)^{+}\!\hskip 0.50003pt\textup{d}x$
$\leq\,$ $\int\limits_{\Omega}\left(|u_{2}(x,t_{1})|^{p-2}u_{2}(x,t_{1})-|u_{1}(x,t_{1})|^{p-2}u_{1}(x,t_{1})\right)^{+}\!\hskip 0.50003pt\textup{d}x\,\,\rightarrow\,\,0,$
when $t_{1}\,\to\,0$. We conclude that $u_{2}\leq u_{1}$ in $\Omega_{T}$. By symmetry $u_{1}\leq u_{2}$.
The same argument proves the corresponding comparison principle for solutions with time derivatives.
Next we state the comparison principle with boundary values
$\psi\in C(\overline{\Omega_{T}})\cap L^{p}(0,T;W^{1,p}(\Omega))$. Observe that now no existence of time derivatives for solutions is assumed.
Theorem 3.1 (Comparison Principle).
Let $u_{1}$ and $u_{2}$ be two weak solutions with boundary values $\psi_{1},\psi_{2}$ as in Definition 2.1. Assume that there is a constant $\varepsilon>0$ such that
(3.3) $\psi_{1}(x,t)\leq\psi_{2}(x,t)-\varepsilon\quad\text{ for all }\quad(x,t)\in\partial_{p}\Omega.$
Then $u_{1}\leq u_{2}$ in $\Omega_{T}$.
We introduce the tools for constructing the required test functions. Fix a convex function $\eta\in C^{\infty}({\mathbb{R}})$ satisfying
$\eta(z):=\begin{cases}0&\text{ when }z\leq 0\\ z-\frac{1}{2}&\text{ when }z\geq 1.\end{cases}$
A key property will be that, upon rescaling, the derivative $\eta^{\prime}$ approximates the Heaviside function and converges to the $\operatorname{sign}$-function. For a parameter $\sigma>0$, we define the two functions
[…]
and also
(3.4) $q_{\sigma,\pm}(z,w):=\int\limits_{w}^{z}\!\eta_{\sigma,\pm}^{\prime}(s\,-\,w)\,b^{\prime}(s)\,\hskip 0.50003pt\textup{d}s.$
Here $\eta_{\sigma,\pm}^{\prime}$ is the derivative of $\eta_{\sigma,\pm}$, and $b^{\prime}(z)=(p-1)|z|^{p-2}$ belonging to $L^{1}_{\text{loc}}({\mathbb{R}})$ is the distributional derivative of $b$. Note that $b^{\prime}(z)>0$ when $z\neq 0$.
Recall that a smooth function $\mathbb{R}$$\to{\mathbb{R}}$ is convex if and only if its first derivative is non-decreasing. Since $\eta$ is a smooth, convex function, so is $\eta_{\sigma,\pm}$ and hence the derivative $\eta_{\sigma,\pm}^{\prime}$ is non-decreasing. This will be crucial in the next lemma.
Lemma 3.2.
For all $z_{1},z_{2},w\in{\mathbb{R}}$ and $\sigma>0$ we have
$q_{\sigma,\pm}(z_{1},w)-q_{\sigma,\pm}(z_{2},w)\,\geq\,\eta_{\sigma,\pm}^{\prime}(z_{2}-w)\,(b(z_{1})-b(z_{2})).$
Proof.
We prove the claim for $\eta_{\sigma,+}$ since the proof for $\eta_{\sigma,-}$ is similar. First assume that $z_{1}\geq z_{2}$ and observe
$q_{\sigma,+}(z_{1},w)-q_{\sigma,+}(z_{2},w)$ $=\,\,\int\limits_{z_{1}}^{z_{2}}\eta_{\sigma,+}^{\prime}(s-w)\,b^{\prime}(s)\,\text{d}s$
$\geq\,\,\eta_{\sigma,+}^{\prime}(z_{2}-w)\int\limits_{z_{1}}^{z_{2}}\,b^{\prime}(s)\,\text{d}s$ $=\,\,\eta_{\sigma,+}^{\prime}(z_{2}-w)\,(b(z_{1})-b(z_{2})).$
In the inequality we used the fact that $\eta_{\sigma,+}^{\prime}$ is non-decreasing. If $z_{1}<z_{2}$, then the estimate reads as
$q_{\sigma,+}(z_{1},w)-q_{\sigma,+}(z_{2},w)$ $=\,\,-\int\limits_{z_{1}}^{z_{2}}\eta_{\sigma,+}^{\prime}(s-w)b^{\prime}(s)\,\text{d}s$
$\,\,\geq\eta_{\sigma,+}^{\prime}(z_{2}-w)\,(b(z_{1})-b(z_{2})).$
This proves the inequality. ∎
We sometimes omit the spatial variable $x$ and write $u(t)=u(x,t)$ and $\nabla u(t)=\nabla u(x,t).$ Finally, let $\phi\in C_{0}^{\infty}({\mathbb{R}})$ be a function satisfying
[…]
For a parameter $\lambda>0$ and a non-negative function $\gamma\in C_{0}^{\infty}(0,T)$, we define the function $\gamma_{\lambda}\in C_{0}^{\infty}({\mathbb{R}}^{2})$ by
(3.5) $\gamma_{\lambda}(t,s):=\frac{1}{\lambda}\,\phi\Bigl(\frac{t-s}{\lambda}\Bigr)\gamma\Bigl(\frac{t+s}{2}\Bigr).$
There is no $x$-dependence here. Thus $\nabla\gamma_{\lambda}=0.$ Later we shall send $\lambda$ to 0.
Lemma 3.3 (Key lemma).
Let $\psi_{1},\psi_{2},u_{1},u_{2}$ and the constant ${\varepsilon}>0$ be as in Theorem 3.1, and let $\gamma\in C_{0}^{\infty}(0,T)$ be non-negative. For all $\lambda$ small enough, all $\sigma>0$, and almost every $s\in(0,T)$, we have
(3.6) $0$ $\geq\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\Bigl\{-q_{\sigma,+}(u_{1}(t),u_{2}(s))\partial_{t}\gamma_{\lambda}(t,s)\Bigr$.
$+$ $\,\Bigl.\gamma_{\lambda}(t,s)\eta_{\sigma,+}^{\prime\prime}(u_{1}(t)\!-\!u_{2}(s))\langle\left|\nabla u_{1}(t)\right|^{p-2}\nabla u_{1}(t),\nabla u_{1}(t)\!-\!\nabla u_{2}(s)\rangle\!\Bigr\}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
Similarly for small enough $\lambda$, all $\sigma>0$, and almost every $t\in(0,T)$, we have
(3.7) $0$ $\geq\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\Bigl\{-q_{\sigma,-}(u_{2}(s),u_{1}(t))\partial_{s}\gamma_{\lambda}(t,s)\Bigr$.
- $\,\Bigl.\gamma_{\lambda}(t,s)\eta_{\sigma,+}^{\prime\prime}(u_{1}(t)\!-\!u_{2}(s))\langle\left|\nabla u_{2}(s)\right|^{p-2}\nabla u_{2}(s),\nabla u_{1}(t)-\nabla u_{2}(s)\rangle\!\Bigr\}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}s.$
Proof.
We first prove (3.6). By the uniform continuity, there is a $\delta>0$ such that
$|\psi_{1}(x,t)-\psi_{2}(x,s)|<\varepsilon/2\quad\text{when}\quad\left|t-s\right|<\delta$
for all $t,s\in[0,T]$ and $x\in\partial\Omega.$ It follows that
$\psi_{1}(x,t)-\psi_{2}(x,s)<-\varepsilon/2.$
By the definition of $\gamma_{\lambda}$, it holds that
$\gamma_{\lambda}(t,s)=0\quad\text{when}\quad\left|t-s\right|\geq\lambda$
for all $t,s\in[0,T]$.
Also for small enough $\lambda>0$ only depending on the support of $\gamma$, it holds that $t\mapsto\gamma_{\lambda}(t,s)\in C_{0}^{\infty}(0,T)$ for all $s>0$. From now on we only consider small enough $\lambda>0$; we see that the choice of $\lambda$ only depends on $\gamma$ and $\delta$.
Recall that $\eta_{\sigma,+}(z)=0$ when $z\leq 0$. We next observe that
$f(x,t):=\eta_{\sigma,+}^{\prime}(u_{1}(x,t)-u_{2}(x,s))\,\gamma_{\lambda}(t,s)$
belongs to $L^{p}(0,T;W_{0}^{1,p}(\Omega))$ for almost every $s\in(0,T)$. We prove that (3.6) holds for such $s$.
For notational convenience, we extend $f(t)=0$ for $t>T$ and $t<0$, and put $u_{1}(t)=\psi_{1}(0)$ for $t<0$. We first consider the Stekloff averages
$f_{h}(x,t):=\frac{1}{h}\int\limits_{t}^{t+h}f(x,\tau)\hskip 0.50003pt\textup{d}\tau.$
Note that $f_{h}\in L^{p}(0,T;W_{0}^{1,p}(\Omega))$ for all small enough $h>0$, and that $f_{h}$ has the time derivative
$\partial_{t}f_{h}(t)=\frac{f(t+h)-f(t)}{h}\in L^{\infty}(\Omega_{T}).$
Hence, we may use $f_{h}$ as a test function in the equation (2.1). This gives
$\int\limits_{0}^{T}\!\!\int\limits_{\Omega}b(u_{1}(t))\partial_{t}f_{h}(t)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\,=\,\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\langle\left|\nabla u_{1}(t)\right|^{p-2}\nabla u_{1}(t),\nabla f_{h}(t)\rangle\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
We also have
$\int\limits_{0}^{T}\!\!\int\limits_{\Omega}b(u_{1}(t))\partial_{t}f_{h}(t)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t$ $=\int\limits_{0}^{T}\!\!\int\limits_{\Omega}b(u_{1}(t))\,\frac{f(t\!+\!h)-f(t)}{h}\,\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t$
$=\int\limits_{0}^{T}\!\!\int\limits_{\Omega}-\,\frac{b(u_{1}(t))-b(u_{1}(t\!-\!h))}{h}f(t)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
The second equality follows from a simple change of variables and the notational conventions made earlier in the proof. By the definition of $f$ and Lemma 3.2, we estimate
$\int\limits_{0}^{T}$ $\!\!\int\limits_{\Omega}\frac{b(u_{1}(t))-b(u_{1}(t-h))}{h}\,f(t)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t$
$=$ $\quad\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\frac{b(u_{1}(t))-b(u_{1}(t-h))}{h}\,\eta_{\sigma,+}^{\prime}(u_{1}(t)\!-\!u_{2}(s))\gamma_{\lambda}(t,s)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t$
$\geq$ $\quad\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\frac{q_{\sigma,+}(u_{1}(t),u_{2}(s))-q_{\sigma,+}(u_{1}(t-h),u_{2}(s))}{h}\,\gamma_{\lambda}(t,s)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t$
$=$ $\quad\int\limits_{0}^{T}\!\!\int\limits_{\Omega}-q_{\sigma,+}(u_{1}(t),u_{2}(s))\,\frac{\gamma_{\lambda}(t+h,s)-\gamma_{\lambda}(t,s)}{h}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
In the last equality we used the fact that $u_{2}(s)$ does not depend on $t$. A combination of the previous three displays gives
$\int\limits_{0}^{T}$ $\!\!\int\limits_{\Omega}-q_{\sigma,+}(u_{1}(t),u_{2}(s))\,\frac{\gamma_{\lambda}(t+h,s)-\gamma_{\lambda}(t,s)}{h}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t$
$\leq$ $\quad-\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\langle\left|\nabla u_{1}(t)\right|^{p-2}\nabla u_{1}(t),\nabla f_{h}(t)\rangle\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
By letting $h\to 0+$, we get
$0$ $\geq\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\Big\{-q_{\sigma,+}(u_{1}(t),u_{2}(s))\,\partial_{t}\gamma_{\lambda}(t,s)$
$\hskip 130.0002pt+\langle\left|\nabla u_{1}(t)\right|^{p-2}\nabla u_{1}(t),\nabla f(t)\rangle\Big\}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t$
$=\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\Big\{-q_{\sigma,+}(u_{1}(t),u_{2}(s))\,\partial_{t}\gamma_{\lambda}(t,s)$
$+$ $\gamma_{\lambda}(t,s)\eta_{\sigma,+}^{\prime\prime}(u_{1}(t)\!-\!u_{2}(s))\langle\left|\nabla u_{1}(t)\right|^{p-2}\nabla u_{1}(t),\nabla u_{1}(t)-\nabla u_{2}(t)\rangle\Big\}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
In the integrand we used the chain rule to compute $\nabla f(t)$. This completes the proof of (3.6).
As for the proof of (3.7), note that $\eta_{\sigma,-}(z)=0$ when $z\geq 0$. Hence, by a similar argument as in the beginning of the proof, we have that
$g(x,s):=\eta_{\sigma,-}(u_{2}(s)-u_{1}(t))\,\gamma_{\lambda}(t,s)$
is contained in $L^{p}(0,T;W_{0}^{1,p}(\Omega))$ for all $\lambda>0$ small enough, all $\sigma>0$ and almost every $t\in(0,T)$. The remainder of the proof is essentially similar as above. ∎
Next we derive fundamental estimates by sending $\sigma\to 0+$ in (3.6) and (3.7).
Proposition 3.4.
Let $\psi_{1},\psi_{2}$ as well as $u_{1},u_{2}$ be as in Theorem 3.1, and let $\gamma\in C_{0}^{\infty}(0,T)$ be a non-negative function. Then for all $\lambda$ small enough we have that
[…]
Proof.
By integrating (3.6) over $s\in(0,T)$ and (3.7) over $t\in(0,T)$, and adding these two integrals together, we get
$0\,\geq\int\limits_{0}^{T}\!\!\!\int\limits_{0}^{T}\!\!\!\int\limits_{\Omega}\!$ $\Bigl\{-q_{\sigma,+}(u_{1}(t),u_{2}(s))\partial_{t}\gamma_{\lambda}(t,s)-q_{\sigma,-}(u_{2}(t),u_{1}(s))\partial_{s}\gamma_{\lambda}(t,s)\Bigr$.
$\Bigl.+\gamma_{\lambda}(t,s)\eta_{\sigma,+}^{\prime\prime}(u_{1}(t)\!-\!u_{2}(s))M(\nabla u_{1}(t),\nabla u_{2}(s))\Bigr\}\,\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\hskip 0.50003pt\textup{d}s,$
where we have written
$M(\nabla u_{1},\nabla u_{2})\,=\,\big\langle\left|\nabla u_{1}\right|^{p-2}\nabla u_{1}-\left|\nabla u_{2}\right|^{p-2}\nabla u_{2},\nabla u_{1}-\nabla u_{2}\big\rangle.$
Note that the term on the second row is non-negative. This is because $\eta_{\sigma,+}^{\prime\prime}\geq 0$ by convexity, $\gamma_{\lambda}\geq 0$ by assumption, and we have the standard inequality $\langle\left|a\right|^{p-2}a-\left|b\right|^{p-2}b,a-b\rangle\geq 0$. It follows that
$\int\limits_{0}^{T}\!\!\!$ $\int\limits_{0}^{T}\!\!\!\int\limits_{\Omega}\!\left(q_{\sigma,+}(u_{1}(t),u_{2}(s))\partial_{t}\gamma_{\lambda}(t,s)+q_{\sigma,-}(u_{2}(t),u_{1}(s))\partial_{s}\gamma_{\lambda}(t,s)\right)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\hskip 0.50003pt\textup{d}s\,$
 $\geq\,0.$
We now prove (3.8) by letting $\sigma\to 0+$. To this end, we verify a suitable convergence. First, note that when $\sigma\to 0+$
$\eta_{\sigma,+}^{\prime}(z)\,\to\,\operatorname{sgn}(z):=\begin{cases}0,&z\leq 0,\\ 1,&z>0,\end{cases}$
uniformly in ${\mathbb{R}}\setminus[-a,a]$ where $a>0$ is fixed. Hence, as $\sigma\to 0+$, we have
$q_{\sigma,+}(z,w)\to(b(z)-b(w))^{+}\quad\text{ pointwise in}\quad{\mathbb{R}}^{2},$
where $(b(z)-b(w))^{+}=\max\{b(z)-b(w),0\}$. By a straightforward computation, we see that
$\left|q_{\sigma,+}(z,w)\right|\leq(b(z)-b(w))^{+}\quad\text{ for all}\quad z,w\in{\mathbb{R}}\,\,\text{and}\,\,\sigma>0.$
There are analogous properties for
$q_{\sigma,-}(z,w)\to(b(w)-b(z))^{+}\quad\text{ and }\quad\left|q_{\sigma,-}(z,w)\right|\leq(b(z)-b(w))^{+}.$
Hence, by letting $\sigma\to 0+$, inequality (3.8) follows from Lebesgue’s dominated convergence theorem. ∎
Next, we let $\lambda\to 0+$ in (3.8) to prove the following estimate.
Proposition 3.5.
Let $\psi_{1},\psi_{2}$ as well as $u_{1},u_{2}$ be as in Theorem 3.1, and let $\gamma\in C_{0}^{\infty}(0,T)$ be non-negative. Then
[…]
Proof.
We consider $\lambda>0$ small enough so that (3.8) holds and $\gamma_{\lambda}\in C_{0}^{\infty}((0,T)\times(0,T))$. Differentiating (3.5), we find
$(\partial_{t}+\partial_{s})\gamma_{\lambda}(t,s)=\frac{1}{\lambda}\phi\Bigl(\frac{t-s}{\lambda}\Bigr)\gamma^{\prime}\Bigl(\frac{t+s}{2}\Bigr).$
Here the problematic terms with the derivative $\phi^{\prime}$ have cancelled. With this choice inequality (3.8) becomes
[…]
By a standard procedure for approximating Dirac’s delta with mollifiers, letting $\lambda\to 0+$ we arrive at
[…]
See Lemma 2.9 on page 76 in [HH15] for details. This completes the proof. ∎
Proof of Theorem 3.1.
Let $0<t_{1}<t_{2}<T$. By replacing $\gamma$ in (3.9) by a suitable cut-off function, approximating the characteristic function of the interval $t_{1},t_{2}]$, we obtain at the limit
$\int\limits_{\Omega}\left(b(u_{1}(t_{2}))-b(u_{2}(t_{2}))\right)^{+}\hskip 0.50003pt\textup{d}x\,\leq\int\limits_{\Omega}\left(b(u_{1}(t_{1}))-b(u_{2}(t_{1}))\right)^{+}\hskip 0.50003pt\textup{d}x.$
Since
$\int\limits_{\Omega}\left(b(u_{1}(t_{2}))-b(u_{2}(t_{2}))\right)^{+}\!\hskip 0.50003pt\textup{d}x\leq$ $\int\limits_{\Omega}\left(b(u_{1}(t_{1}))-b(\psi_{1}(t_{1})\right)^{+}\!\hskip 0.50003pt\textup{d}x$
$+\int\limits_{\Omega}\left(b(\psi_{1}(t_{1}))-b(\psi_{2}(t_{1})\right)^{+}\!\hskip 0.50003pt\textup{d}x$
$+\int\limits_{\Omega}\left(b(u_{2}(t_{1}))-b(\psi_{2}(t_{1}))\right)^{+}\!\hskip 0.50003pt\textup{d}x.$
The first and the third terms on the right-hand side converge to zero as $t_{1}\to 0$, and the second term is zero by (3.3). Thus
$\int\limits_{\Omega}\bigl(b(u_{1}(t_{2}))-b(u_{2}(t_{2}))\bigr)^{+}\hskip 0.50003pt\textup{d}x\,\,\leq\,\,0,$
and from which we conclude that $u_{1}(x,t_{2})\leq u_{2}(x,t_{2}).$ Since $t_{2}$ is arbitrary, this completes the proof of the comparison principle. ∎
4. Proof of the uniqueness
In this section we prove the main result of the paper, uniqueness with sign-changing $C^{2}$ boundary data. Our proof relies on properties of suitably regular auxiliary solutions. Given $\psi\in C^{2}(\overline{\Omega_{T}})$, we say that a weak solution $u\in C(\overline{\Omega_{T}})$ with boundary values $\psi$ is $t$-regular if the Sobolev time derivative $\partial_{t}(\left|u\right|^{p-2}u)$ exists and
$\int\limits_{0}^{T}\!\!\int\limits_{\Omega}(\left|\partial_{t}b(u)\right|^{2}+\left|\nabla u\right|^{p})\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\,<\,\infty.$
We now note that there exists one and only one such $t$-regular solution, see [LPS25a, Pages 4-6] for details. They are unique at least within their own regularity class. The next theorem guarantees that there are no other weak solutions
Theorem 4.1.
A solution with boundary values $\psi\in C^{2}(\overline{\Omega_{T}})$ is unique.
Proof.
Let $u$ be an arbitrary solution with the given boundary values $\psi$. According to [LPS25a, Theorem 5], there always exists a $t$-regular solution with the boundary values $\psi\in C^{2}(\overline{\Omega_{T}})$. Moreover, it is well-known that such a $t$-regular solution is unique, see for example [LPS25a, Corollary 3], although the existence of less regular solutions is not yet excluded. Given a constant $\varepsilon>0$, there exists unique $t$-regular solutions
$\begin{cases}u_{+\varepsilon}\quad\text{with boundary values}\quad\psi+{\varepsilon}\\ \,u_{0}\quad\,\,\text{with boundary values}\quad\psi+0\\ u_{-\varepsilon}\quad\text{with boundary values}\quad\psi-{\varepsilon}.\end{cases}$
On the other hand, by the comparison principle in Theorem 3.1, we have for almost every $(x,t)\in\Omega\times(0,T)$ that
(4.10) $u_{-\varepsilon}(x,t)\leq u(x,t)\leq u_{+\varepsilon}(x,t).$
It is essential that the above comparison principle does not require $u$ to be $t$-regular. We claim that $u=u_{0}$.
Finally, since $\psi\in C^{1}(\overline{\Omega_{T}})$ and $\|\partial_{t}\nabla\psi\|_{L^{p}(\Omega_{T})}<\infty$ by our assumption, we may use [LPS25a, Theorem 5 and Corollary 3] to conclude that, upon passing to a subsequence if necessary,
$u_{\pm\varepsilon}\,\to\,u_{0}\quad\text{in}\quad L^{p}(0,T;W^{1,p}(\Omega))\quad\text{as}\quad\varepsilon\to 0.$
The point is that both sequences have the same limit, viz. $u_{0}$. For a subsequence the convergences hold pointwise for almost every $(x,t)\in\Omega\times(0,T)$. This together with (4.10) implies that $u=u_{0}$, and thus the proof of the uniqueness is complete. ∎
The proof has an interesting consequence. Since we obtained $u=u_{0}$, we can read off the following result from the estimate in [LPS25a, Theorem 5].
Corollary 4.2.
A weak solution with $C^{2}$ boundary data possesses the Sobolev derivatives
$\partial_{t}(|u|^{\frac{p-2}{2}}u)\,\in L^{2}(\Omega_{T}),\qquad p\geq 2$
$\partial_{t}(|u|^{p-2}u)\,\in\,L^{p}(\Omega_{T}),\qquad 1<p<2.$